Calculate the mass of solution in a 6 ppm solution of a drug if the mass of the solute is 0.050 mg.
1
Expert's answer
2020-11-09T11:51:50-0500

We know that: 6 ppm  = 6 mg/l

6 mg/l = 0.050/x

Volume of solute (x) = 0.050/6 = 0.008 l

Assuming density of solute is 1 g/ml = 1 kg/l,

actual mass of the solute is:

m (solute) = 0.008 kg = 8 g = 8000 mg

mass of solution = mass of solute + mass of solvent = 0.050 mg + 8000 = 8000.05 mg